Find dy/dx by implicit differentiation. 4x2 + 7xy - y2 = 3 y =
Answer #1

Given, 4x² + 7xy-y²= 3 - We Knowe that I alten) = nacht by product rule d lu.v) dre =u du tre + v. du du ☺ worito k 0 Ju Diff